If m = 2, find the value of:
3m2   – 2m – 7...............

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that m = 2
Then, substitute the value of p in the question
= (3 × 22)–(2 × 2) - 7
= (3 × 4)–(4) - 7
= 12–4 -7
= 12–11
= 1

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Given : m = 2
Substituting value of m in the equation,
3m - 2m -7
= 3×2 - 2×2 - 7
= 6-4-7 
= 6-11
= -5
 
Hope it helps!
Thank you ✌️